OpenStudy (anonymous):

please help 1 medal and fan Solve 7/4=3/x. Round to the nearest tenth. a. 1.7 b. 3.7 c. 2.5 d. 1.2

OpenStudy (anonymous):

okay so 7/4=3/x can be rewritten as x/4=3/7 rearrange the formula for x=(3*4)/7 then take to one decimal place (for tenths) and looking at the second decimal point either round up or down accordingly

OpenStudy (anonymous):

1.7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

correct! no problem :)